Write the slope-intercept equation for the line that passes through (2, -4) and has slope -7/2. Please show all of your work.
Answer by mananth(16949)

m=-3.5

Plug value of the slope and point (2, -4) in
Y =mx+b
-4.00=-7+b
b= -4+7
b=3
So the equation will be
Y =-3.5x+3
m.ananth@hotmail.ca